Jump to main content.


Drake Chemical

EPA Weekly Situation Reports for February 1999

U. S. ARMY CORPS OF ENGINEERS
DRAKE CHEMICAL SUPERFUND SITE

182 MYRTLE STREET
LOCK HAVEN, PA 17745
PHONE 717-748-4505
FAX 717-748-7549

Wednesday, February 24, 1999

FOR DATES: From 7:00 a.m. Wednesday, February 17, 1999 to 7:00 a.m. Wednesday, February 24, 1999.

The Thermal Destruction Facility (TDF) was on-line for 165.7 hours and processed 4289.7 tons of contaminated soil during the time frame of this report. The TDF was on-line 98.6 % of the time with an average hourly feed rate of 25.9 tons per hour while on-line. The contractor, OHM, continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated approximately 257,267 tons of contaminated soil, which amounts to approximately 89.0 % of the overall contaminated soil on-site. The overall percentage complete is based on a new the site survey done last week. The survey used a smaller grid profile than previous surveys and the estimated quantity remaining to be treated increased by 5000 tons.

Hazardous waste feed to the incinerator was stopped briefly on Tuesday, February 23, 1999 in order to repair the quench tower screw conveyor. Repairs to the screw conveyor consisted of replacing several bolts that connected a coupling at the screws drive shaft.

Site activities during the time frame of this report included TDF operations and maintenance, wastewater management, equipment decontamination, wood shredding, feed preparation, debris handling, debris separation, lime addition, ash handling, ash backfilling, dust control, excavation surveys, dry decontamination and backfilling of rock and concrete debris. The contractor continued to perform off-site and perimeter air monitoring (PAM) activities.

Backfill operations continued adjacent to the shoring systems of the ash stabilization pad. Imported clean shot rock was used to backfill the 2-foot rock bottom in these areas. The sheet piling on the exclusion zone side of the utility corridor was cut below existing grade in order to ease the future mass grading of clean ash, once the incineration phase is complete.

During the time frame of this report, OHM began implementing the site close out wastewater management plan. The frac tanks that will hold future storm water and decontamination water, in lieu of the sediment basin, were staged near their final position. In addition, the secondary containment piping that will house the future water management piping was installed across the exclusion zone to the WWTP.

Excavation activities were continued near the southern end of the exclusion zone and adjacent to the ash stabilization pad shoring system. Excavated soils consisted primarily of existing fill and extremely wet lagoon soil, both were underlain by naturally occurring soil. During the time frame of this report, no intact drums or drum carcasses were unearthed during excavation activities.

During the time frame of this report, two roll-off containers of PPE and dry decontaminated miscellaneous metal were shipped off-site as non-hazardous waste to a hazardous waste landfill. One roll-off containing asbestos piping was shipped off-site to a hazardous waste landfill. One roll-off of decontaminated scrap steel was shipped off-site for recycling.

Related site activities included:

* A weekly progress meeting was held on Tuesday, February 23, 1999, with representatives from the USEPA, PADEP, USACE, and OHM in attendance.

* A senior partnering meeting and the monthly MIS meeting were held on Tuesday, February 23, 1999, with representatives from the USEPA, PADEP, USACE, and OHM in attendance.

Major activities for the next week include: Operation of the TDF, soil feed preparation, excavation, drum handling and drum bulking (if encountered), lime treatment of soil, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning, rock backfilling, and the implementation of the project close out plan.

PREPARED BY: Steve Gibson for William Werntges (USACE)


Wednesday, February 17, 1999

FOR DATES: From 7:00 a.m. Wednesday, February 10, 1999 to 7:00 a.m. Wednesday, February 17, 1999.

The Thermal Destruction Facility (TDF) was on-line for 143.1 hours and processed 4694.3 tons of contaminated soil during the time frame of this report. The TDF was on-line 85.2 % of the time with an average hourly feed rate of 32.8 tons per hour while on-line. The contractor, OHM, continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated approximately 254,690.2 tons of contaminated soil, which amounts to approximately 89.7 % of the overall contaminated soil on-site.

Hazardous waste feed to the incinerator was stopped on Friday, February 12, 1999 at 0800 hours in order to perform scheduled maintenance and monthly TDF calibrations. Maintenance activities included the application of hard face and cutting teeth to the dual feed screws. Monthly TDF calibrations and monthly AWFSO checks were performed to test the incinerators automatic shut down systems. As part of these interlock checks, a planned Thermal Relief Valve (TRV) opening was initiated at 1421 hours. Prior to the TRV opening, feed to the incinerator had been off-line for approximately 6.5 hours and the proper local authorities were notified. Hazardous waste feed was resumed at 0441 hours on Saturday, February 13, 1999, after repairs to a sensor on the emergency generator had been made.

During the time frame of this report, the incinerator was also off-line on Saturday evening, February 13, 1999, in order to reduce the specific gravity of the scrubber water. The acid gas scrubber water's specific gravity had raised slightly above normal operating range, and scrubber water blow down rate was increased temporarily until the readings taken were within normal operating ranges. Hazardous waste feed was off-line for approximately 3.5 hours.

Site activities during the time frame of this report included TDF operations and maintenance, wastewater management, equipment decontamination, wood shredding, feed preparation, debris handling, debris separation, lime addition, ash handling, ash backfilling, dust control, excavation surveys, dry decontamination and backfilling of rock and concrete debris. The monthly survey of the exclusion zone was performed to determine the quantity of material remaining in the exclusion zone awaiting thermal treatment. The contractor continued to perform off-site and perimeter air monitoring (PAM) activities.

Backfill operation are continuing adjacent to the shoring systems at the corridor road and ash stabilization pad. Imported clean shot rock is being used along with the cleaned rock at the 2' bottom. Once the rock is graded and confirmed to be above elevation 547, clean bottom ash is placed and compacted along these areas.

Excavation activities continue near the southern end of the exclusion zone and adjacent to the ash stabilization pad shoring system. Excavated soils consisted primarily of existing fill and site sludge, which were underlain by natural soil. During the time frame of this report, no intact drums or drum carcasses were unearthed during excavation activities.

Two roll-off containers of PPE and dry decontaminated drum carcasses were shipped off-site as non-hazardous waste to a hazardous waste landfill. The activated carbon, which is used as a treatment component of site water in the site WWTP, was transferred from the carbon vessels to a tanker truck for shipment off-site for regeneration. The carbon vessels were then refilled with fresh activated carbon.

Related site activities included:

* A weekly progress meeting was held on Tuesday, February 16, 1999, with representatives from the USEPA, PADEP, USACE, and OHM in attendance.

* A site tour was given to Science Teachers from the Williamsport Area School District high school and middle schools.

* A site tour was given to a group of environmental law students from Penn State University.

Major activities for the next week include: Operation of the TDF, soil feed preparation, excavation, drum handling and drum bulking (if encountered), lime treatment of soil, off-site and perimeter air monitoring activities, WWTP operations, sheet pile reinforcing, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning, and rock backfilling.

PREPARED BY: Steve Gibson for William Werntges (USACE)


Wednesday, February 10, 1999

FOR DATES: From 7:00 a.m. Wednesday, February 3, 1999 to 7:00 a.m. Wednesday, February 10, 1999.

The Thermal Destruction Facility (TDF) was on-line for 166.4 hours and processed 5786.9 tons of contaminated soil during the time frame of this report. The TDF was on-line 99.1 % of the time with an average hourly feed rate of 34.8 tons per hour while on-line. The contractor, OHM, continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated approximately 248,734.4 tons of contaminated soil, which amounts to approximately 87.6 % of the overall contaminated soil on-site.

Hazardous waste feed to the incinerator was interrupted briefly on February 3, 1999 in order to replace a roller on the weighbelt conveyance system. Miscellaneous down time to the TDF was attributed to performance of the daily calibrations to the weighbelt and to clear excess soil from the dual feed screws.

Site activities during the time frame of this report included TDF operations, wastewater management, wood shredding, feed preparation, debris handling, debris separation, lime treatment of soils, ash handling, ash backfilling, dust control to prevent fugitive emissions, site surveys, dry decontamination and backfilling of rock and concrete debris. OHM continued to decontaminate several large pieces of equipment for removal from the exclusion zone. The equipment will either be shipped off-site or used for spare parts. The contractor continued to perform off-site and perimeter air monitoring (PAM) activities. The monthly sampling event for TSP/metals and semi-volatile compounds was performed. Nestlerode Construction completed the installation of the excavation support shoring system and demobilized from the site.

Excavation activities were continued near the southern end of the exclusion zone and adjacent to the ash stabilization pad sheet steel. During the time frame of this report, no intact drums or drum carcasses were unearthed during excavation activities.

Two roll-off containers of non-hazardous PPE and lime bags were shipped off-site to a hazardous waste landfill on February 5, 1999. The contractor sampled the content of four drums during the time frame of this report. The samples of drum contents were shipped off-site for fingerprint analyses and compatibility testing.

The contents of seventy-two (72) recently excavated drums were combined in a lined roll-off as part of the drum bulking activities. Prior to bulking activities, the drum contents had been sampled for compatibility, as indicated in the previous Weekly Situation Reports. Once all of the compatible drum contents are combined, the contents will be homogenized and sampled to determine how the contents will be manifested for shipment to an offsite disposal facility.

Related site activities included a weekly progress meeting that was held on Tuesday, February 9, 1999, with representatives from the USEPA, PADEP, USACE, and OHM in attendance.

Major activities for the next week include: Operation of the TDF, soil feed preparation, excavation, drum handling, drum sampling, drum bulking, lime treatment of soil, off-site and perimeter air monitoring activities, WWTP operations, sheet pile reinforcing, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning, and rock backfilling.

On Friday, February 12, 1999 feed to the incinerator will be stopped and the TDF will be placed in hot hold in order to perform preventative maintenance and to perform the monthly TDF calibrations. Maintenance activities will include the application of hard face and cutting teeth to the dual feed screws. The monthly interlock checks, which test the incinerators automatic shut down systems, are scheduled to be performed during the down time. As part of the interlock checks, a planned Thermal Relief Valve (TRV) opening will be initiated. Prior to opening the TRV, the incinerator will be purged of hazardous materials and the proper local authorities will be notified. The incinerator is scheduled to be down for approximately 10 hours.

PREPARED BY: Steve Gibson for Chad Thompson (USACE)


Wednesday, February 3, 1999

FOR DATES: From 7:00 a.m. Wednesday, January 27, 1999 to 7:00 a.m. Wednesday, February 3, 1999.

The Thermal Destruction Facility (TDF) was on-line for 167.0 hours and processed 5703.9 tons of contaminated soil during the time frame of this report. The TDF was on-line 99.4 % of the time with an average hourly feed rate of 34.1 tons per hour while on-line. The contractor, OHM, continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated approximately 242,428 tons of contaminated soil, which amounts to approximately 85.5 % of the overall contaminated soil on-site.

Activities during the time frame of this report included TDF operations, wastewater management, wood shredding, feed preparation, debris handling, debris separation, lime treatment of soils, ash handling, ash backfilling, dust control to prevent fugitive emissions, dry decontamination and backfilling of rock and concrete debris. Nestlerode Construction continues to install the whaler support system in order to stiffen the sheet piling in the exclusion zone. As part of OHM's demobilization plan, several large pieces of equipment were decontaminated and removed from the exclusion zone to either be shipped off-site or used for spare parts.

The contractor continued to perform off-site and perimeter air monitoring (PAM) activities. At 0722 hours on Saturday, January 30, 1999, an NMOC alarm was recorded at PAM station 2. Investigation indicated that the alarm was the result of a short circuit in the #2 heat trace line and not an actual detection. While repairs were being made to the system, site air was monitored with hand held devices.

Excavation activities were continued along the southern end of the exclusion zone adjacent to the utility corridor and the ash stabilization pad, the area where Nestlerode Construction had recently installed the H-pile support system. The excavation crew uncovered discarded drums, drum carcasses and demolition debris in the general vicinity of the former debris separation building and rock wash pad. Additional drum carcasses and 4 intact drums were unearthed and separated during this reporting period. The initial Health and Safety scan of each intact drum indicated that the drums contained non-volatile product, and the drums were wrapped in polyethylene for future sampling and overpacking.

Two roll-off containers of non-hazardous PPE and metal debris were shipped off-site to a hazardous waste landfill on February 2, 1999. No drum sampling was performed during the time frame of this report.

Related site activities included:

* A weekly progress meeting was held on Tuesday, February 2, 1999, with representatives from the USEPA, PADEP, USACE, and OHM in attendance.

* Kurt Davis, A.I.R. Group representative, was given a site tour on Wednesday, January 27, 1999.

* Bill Smedley, A.I.R. Group, and Trina Allen, TAG Representative, were given a site tour on February 2, 1999.

Major activities for the next week include: Operation of the TDF, soil feed preparation, excavation, drum handling, drum sampling, lime treatment of soil, off-site and perimeter air monitoring activities, WWTP operations, sheet pile reinforcing, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning, and rock backfilling.

PREPARED BY: Steve Gibson for William Werntges(USACE)

Region 3 | Mid-Atlantic Cleanup | Mid-Atlantic Superfund |EPA Home | EPA Superfund Homepage


Local Navigation


Jump to main content.